Solve the following equations: 
 5t + 28 = 10 ….........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

By transposing 28 from LHS to RHS it becomes -28
Then,
= 5t = 10–28
= 5t = - 18
Now,
Divide both side by 5,
= 5t/5= -18/5
= t = -18/5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
5t + 28 = 10
5t = 10 - 28
5t = -18
t = -18/5
 
Hope it helps!
Thank you ✌️